# Regularization in Data-driven Predictive Control: A Convex Relaxation Perspective

### v1

@@ 2.2 Direct data-driven predictive control
We here review the notion of persistent excitation (PE) to ensure the offline data is sufficiently rich.
@@
The following result from [15], commonly known as the Fundamental lemma, forms the foundation of many recent results of direct data-driven predictive control.
Lemma 1 (Fundamental Lemma [15]).
@@
The fundamental lemma gives a parameterization of all finite-dimensional input/output trajectories of (1) using only one offline trajectory $\{u_{\textnormal{d}},y_{\textnormal{d}}\}$. In particular, the image of the Hankel matrix in (3) is the same as the set of all system trajectories of length $L$. This serves as the foundation of many recent direct data-driven analysis and control methods; see [4, 40, 41] for excellent surveys.
With the fundamental lemma, we can use (3) to build a predictive model to replace (2a)-(2c). This is the basic idea in DeePC [16]. In particular, the Hankel matrix formed by the offline data in (3) is partitioned as
$\begin{bmatrix}U_{\textrm{P}}\\ U_{\textrm{F}}\end{bmatrix}:=\mathcal{H}_{L}(u_{\textrm{d}}),\quad\begin{bmatrix}Y_{\textrm{P}}\\ Y_{\textrm{F}}\end{bmatrix}:=\mathcal{H}_{L}(y_{\textrm{d}}),$ (4)
where $U_{\textnormal{P}}$ and $U_{\textnormal{F}}$ consist the first $T_{\textnormal{ini}}$ rows and the last $N$ rows of $\mathcal{H}_{L}(u_{\textnormal{d}})$, respectively (similarly for $Y_{\textnormal{P}}$ and $Y_{\textnormal{F}}$). The length of offline data satisfies $L\!=\!T_{\textnormal{ini}}\!+\!N$. We denote the most recent past input trajectory of length $T_{\textnormal{ini}}$ and the future input trajectory of length $N$, respectively, as
$u_{\textnormal{ini}}$ $=\textnormal{col}(u(t-T_{\textrm{ini}}),u(t-T_{\textrm{ini}}+1),\ldots,u(t-1)),$
$\bar{u}$ $=\textnormal{col}(\bar{u}(t),\bar{u}(t+1),\ldots,\bar{u}(t+N-1)),$
(similarly for $y_{\textrm{ini}},\bar{y}$). Then, Lemma 1 ensures that the sequence $\textnormal{col}(u_{\textrm{ini}},y_{\textrm{ini}},\bar{u},\bar{y})$ is a valid trajectory of (1) if and only if there exists $g\in\mathbb{R}^{T-T_{\textrm{ini}}-N+1}$ such that
$H_{\mathrm{d}}g=\textnormal{col}(u_{\textrm{ini}},y_{\textrm{ini}},\bar{u},\bar{y}),$ (5a)
@@
If $T_{\mathrm{ini}}$ is larger or equal to the lag of (1), $\bar{y}$ is unique given any $u_{\textrm{ini}}$, $y_{\textrm{ini}}$ and $\bar{u}$ in (5).
The basic DeePC [16] utilizes (5) as the data-driven representation of (2a)-(2c) and reformulate problem (2) as
$\min_{g,\bar{u},\bar{y}}$ $\sum_{k=t}^{t+N-1}\big{(}\|\bar{y}(k)\|_{Q}^{2}+\|\bar{u}(k)\|_{R}^{2}\big{)}$ (6)
@@
where we slightly abuse the notation and use $\bar{u}\in\mathcal{U},\bar{y}\in\mathcal{Y}$ to denote input/output constraints (2d).
@@
For LTI systems with noise-free data, model-based control (2) and DeePC (6) are equivalent (see [16, Theorem 5.1]), thanks to the fundamental lemma. In this case, the Hankel matrix $H_{\mathrm{d}}$ in (5b), also referred to as the trajectory matrix (since each of its columns represents a valid system trajectory), serves as a non-parametric model.
@@
However, data collected from practical systems is rarely noise-free. In particular, the offline data $u_{\mathrm{d}},y_{\mathrm{d}}$ and the resulting trajectory library $H_{\mathrm{d}}$ in (5b) may be corrupted by 1) “variance” noises that enter the process dynamics and output measurement, 2) and “bias” errors induced by nonlinear dynamics beyond LTI [1]. To address these issues, various regularization and data preprocessing techniques have been proposed to extend the basic DeePC (6). These include $l_{1}$$l_{2}$ regularization [1], $\gamma$-DDPC [31], low-rank approximation [36], singular-value decomposition [26]. While some works [26, 1, 42, 31] have explored the relationship among different schemes, most of them are carried out on a case-by-case basis. One notable exception is [1], which employs a principled bi-level optimization framework to investigate the interplay between direct and indirect data-driven control.
@@
We here adopt the same bi-level optimization principle as [1] to systematically explore the effects of regularization in direct Data-driven Predictive Control (DDPC). In particular, we consider a nested bi-level DDPC,
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2},$
$\mathrm{subject~to}\quad$ ${H}^{\star}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (7a)
$\textrm{where}\quad{H}^{\star}\in\arg\min_{{H}\in\mathcal{S}}\;J_{\mathrm{id}}({H},H_{\mathrm{d}}),$ (7b)
@@
where we process the offline data $H_{\mathrm{d}}$ (see (5)) in the inner identification (7b) before using it for the outer online control (7a). This bi-level formulation, often referred to as indirect DDPC, is modular and consists of two well-separated components: an inner identification layer and an outer model-based predictive control layer. Before discussing its connections and distinctions with existing schemes, let us further clarify the notation in (7). In the outer problem, the term $\|y\|_{Q}^{2}+\|u\|_{R}^{2}$ denotes the usual one $\sum_{k=t}^{t+N-1}\big{(}\|y(k)\|_{Q}^{2}+\|u(k)\|_{R}^{2}\big{)}$, and we have introduced a slack variable $\sigma_{y}$ with constraint $\Gamma$ and a regularization term $\|\sigma_{y}\|_{2}^{2}$ with $\lambda_{y}>0$. This ensures constraint feasibility despite noise, as used in [4, 18]. In the inner problem, $J_{\mathrm{id}}(\cdot,\cdot)$ denotes a suitable identification loss, and the constraint ${H}\in\mathcal{S}$ enforces some prior data structures from LTI systems. We will clarify our choice of ${H}\in\mathcal{S}$ in Section 3.
Remark 1 (Comparison with [1]).
@@
The bi-level DDPC (7) is inspired by the recent work [1], which argues that “direct and regularized data-driven control can be derived as a convex relaxation of the indirect approach”. While adopting the same general framework, our work introduces two key distinctions. Conceptually, the bi-level DDPC (7) explicitly emphasizes the use of a non-parametric trajectory matrix as the predictive model, whereas [1] primarily considers a more general and abstract behavioral LTI setting in the inner problem. Technically, the analysis in [1] relies on arguments that require additional assumptions or clarifications, and some claims may not hold without these refinements. Also, the exact penalty arguments in [1] rely on [43, Proposition 2.4.3], which requires adaptation in the DDPC context. Inspired by the concept of partial calmness, our proof is elementary, transparent, and self-contained. We will clarify these points in more detail in Sections 3 and 4.
@@ 3.1 Behavior-based identification for predictive control
From the offline noisy data $H_{\mathrm{d}}$, we aim to get a new trajectory library $H^{\star}$ for predictive control
${H}^{\star}:=\textnormal{col}({U}_{\textnormal{P}}^{\star},{Y}_{\textnormal{P}}^{\star},{U}_{\textnormal{F}}^{\star},{Y}_{\textnormal{F}}^{\star}),$
where each column of $H^{\star}$ is a “purified” trajectory of the system. Let $\mathcal{B}_{L}$ (where $L=T_{\textnormal{ini}}+N$) denotes the space of all possible length-$L$ trajectories of an LTI system, i.e.,
@@
$H^{\star}=\operatorname*{argmin}_{H}$ $J_{\mathrm{id}}({H},H_{\mathrm{d}})=\|H-H_{\mathrm{d}}\|_{F}^{2}$ (8)
$\mathrm{subject~to}$ $\mathrm{Im}\,{H}\subseteq\mathcal{B}_{L}.$
If the offline data $H_{\mathrm{d}}$ is noise-free and satisfies the persistent excitation, the fundamental lemma ensures that $\mathrm{Im}\,{H}_{\mathrm{d}}=\mathcal{B}_{L}$. In this case, the optimal solution to (8) is $H^{\star}={H}_{\mathrm{d}}$. In general, the constraint $\mathrm{Im}\,{H}\subseteq\mathcal{B}_{L}$ is too abstract and not tractable. For an explicit expression, we have several necessary conditions.
  • First, the trajectory matrix $H$ should be of low rank, i.e., we have $\mathrm{rank}\,H^{\star}\leq mL+n$.
@@
    ${Y}_{\textnormal{F}}^{\star}=K\ \textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}}),$ (9)
@@
  • Third, the trajectory matrix $H$ should satisfy causality, i.e., the coefficient matrix $K$ in (9) should have a block partition $K=\begin{bmatrix}K_{\mathrm{p}}\;K_{\mathrm{f}}\end{bmatrix},\ K_{\mathrm{f}}\in\mathcal{L}$, where $K_{\mathrm{f}}\in\mathcal{L}$ encodes a block-low triangle structure for causality.
  • Finally, we may also impose a Hankel structure on $H^{\star}\in\mathcal{H}$ if each column is a one-step shifted trajectory.
@@
$({H}^{\star},K^{\star})\in\operatorname*{argmin}_{\tilde{H},K}\quad$ $\|\tilde{H}-H_{\mathrm{d}}\|_{F}^{2}$
@@
We thus consider (10) as our inner identification in (7b). Note that the requirements of linearity (10b) and causality (10c) are typically used in causal SPC [35, 17], and the requirement of low rank (10a) and Hankel structure (10d) are often employed in low-rank approximation [36]. When the offline data $H_{\mathrm{d}}$ in (5) is collected from an LTI system (1) with noise-free data, the unique optimal solution to (10) is trivially ${H}^{*}=H_{\mathrm{d}}$. If $H_{\mathrm{d}}$ contains “variance” noises and/or “biases” errors, the identification problem (10) becomes challenging to solve.
@@ 3.2 A direct approach via penalty methods
With the inner identification (10), the bi-level formulation (7) is an indirect DDPC. Using penalty methods, the bi-level structure can be transformed into a single-level problem. This is a standard idea in bi-level optimization [44] and has been recently used to analyze direct data-driven control in [1]. We adopt a similar strategy here, but also highlight a subtle point that leads to some conceptual inaccuracies in [1].
@@
$\left\{\begin{aligned} p(H)=0,&\quad\textrm{if}\ {H}\in\mathcal{C}_{\mathrm{opt}},\\ p(H)>0,&\quad\textrm{if}\ {H}\notin\mathcal{C}_{\mathrm{opt}}.\end{aligned}\right$. (12)
@@
Under a mild assumption (see e.g., [45, Theorem 6.6]), we can ensure that as the penalty parameter $\lambda\to\infty$, the optimal solution of (13) converges to an optimal solution of (11), i.e., the original bi-level problem (7). If we use a small penalty parameter $\lambda$, the single-level formulation (13) becomes a relaxation of (7), i.e., the optimal value of (13) is smaller than that in (7).
The reasoning above critically relies on the conceptual optimality constraint $H\in\mathcal{C}_{\mathrm{opt}}$, associated with the inner problem (7b). Dropping any constraint in (7b) results in a different optimality set, meaning that the corresponding penalized single-level problem may no longer offer a valid relaxation. This subtle but important point may be easily overlooked; for example, some relaxation statements in [1] are not fully accurate without further clarification. We consider a simple example below.
@@
It can be verified that when $\lambda\geq\frac{1}{4}$ (see Figure 2 for illustration), (15) has the same optimal solution as (14). If $\lambda<\frac{1}{4}$, the optimal value of (15) is a lower bound of (14). If we relax the constraint (14c) as $y\leq 4$, then the inner problem has a different optimal solution $y^{*}=0$. In this case, the outer problem in (14a) becomes infeasible. This illustrates that dropping constraints from the inner problem does not necessarily reduce the optimal value of the outer problem. This nuance was overlooked in [1].

### v2

@@ 2.2 Direct data-driven predictive control
We here review the notion of persistent excitation (PE), which ensures the offline data is sufficiently rich.
@@
The following result from [willems2005note], commonly known as the fundamental lemma, forms the foundation of many recent results of direct data-driven predictive control.
Lemma 1 (Fundamental Lemma [willems2005note]).
@@
The fundamental lemma gives a parameterization of all finite-dimensional input/output trajectories of (1) using only one offline trajectory $\{u_{\textnormal{d}},y_{\textnormal{d}}\}$. In particular, the image of the Hankel matrix in (3) is the same as the set of all system trajectories of length $L$.
With the fundamental lemma, we can use (3) to build a predictive model to replace (2a)-(2c). This is the basic idea in DeePC [coulson2019data]. In particular, we consider the input/output sequence with length $L$ satisfying $L=T_{\textnormal{ini}}+N$. The Hankel matrix formed by the offline data in (3) is partitioned as
$\begin{bmatrix}U_{\textrm{P}}\\ U_{\textrm{F}}\end{bmatrix}:=\mathcal{H}_{L}(u_{\textrm{d}}),\quad\begin{bmatrix}Y_{\textrm{P}}\\ Y_{\textrm{F}}\end{bmatrix}:=\mathcal{H}_{L}(y_{\textrm{d}}),\vskip-5.69054pt$ (4)
where $U_{\textnormal{P}}$ and $U_{\textnormal{F}}$ consist of the first $T_{\textnormal{ini}}$ rows and the last $N$ rows of $\mathcal{H}_{L}(u_{\textnormal{d}})$, respectively (similarly for $Y_{\textnormal{P}}$ and $Y_{\textnormal{F}}$). We denote the most recent past input trajectory of length $T_{\textnormal{ini}}$ and the future input trajectory of length $N$, respectively, as
$u_{\textnormal{ini}}$ $=\textnormal{col}(\bar{u}(t-T_{\textrm{ini}}),\bar{u}(t-T_{\textrm{ini}}+1),\ldots,\bar{u}(t-1)),$
$u$ $=\textnormal{col}(u(t),u(t+1),\ldots,u(t+N-1)),$
(similarly for $y_{\textrm{ini}},y$). Then, Lemma 1 ensures that the sequence $\textnormal{col}(u_{\textrm{ini}},y_{\textrm{ini}},u,y)$ is a valid trajectory of (1) if and only if there exists $g\in\mathbb{R}^{T-T_{\textrm{ini}}-N+1}$ such that
$H_{\mathrm{d}}g=\textnormal{col}(u_{\textrm{ini}},y_{\textrm{ini}},u,y),$ (5a)
@@
If $T_{\mathrm{ini}}$ is larger or equal to the lag of (1), $y$ is unique given any $u_{\textrm{ini}}$, $y_{\textrm{ini}}$ and $u$ in (5).
The basic DeePC [coulson2019data] utilizes (5) as the data-driven representation of (2a)-(2c) and reformulate problem (2) as
$\min_{g,u,y}$ $\sum_{k=t}^{t+N-1}\big(\|y(k)\|_{Q}^{2}+\|u(k)\|_{R}^{2}\big)$ (6)
@@
where we slightly abuse the notation and use $u\in\mathcal{U},y\in\mathcal{Y}$ to denote input/output constraints (2d).
@@
For LTI systems with noise-free data, model-based control (2) and DeePC (6) are equivalent (see [coulson2019data, Theorem 5.1]), thanks to the fundamental lemma. In this case, the Hankel matrix $H_{\mathrm{d}}$ in (5b), also referred to as the trajectory matrix (since each of its columns represents a valid system trajectory), serves as a non-parametric model.
@@
However, data collected from practical systems is rarely noise-free. In particular, the offline data $u_{\mathrm{d}},y_{\mathrm{d}}$ and the resulting trajectory library $H_{\mathrm{d}}$ in (5b) may be corrupted by 1) “variance” noises that enter the process dynamics and output measurement, 2) and “bias” errors induced by nonlinear dynamics beyond LTI [dorfler2022bridging]. To address these issues, various regularization and data preprocessing techniques have been proposed to extend the basic DeePC (6). These include $l_{1}$$l_{2}$ regularization [dorfler2022bridging], $\gamma$-DDPC [sader2023causality], low-rank approximation [markovsky2016missing], singular-value decomposition [zhang2023dimension]. While some works [zhang2023dimension, dorfler2022bridging, mattsson2024equivalence, sader2023causality] have explored the relationship among different schemes, most of them are carried out on a case-by-case basis. One notable exception is [dorfler2022bridging], which employs a principled bi-level optimization framework to investigate the interplay between direct and indirect data-driven control.
@@
We here adopt the same bi-level optimization principle as [dorfler2022bridging] to systematically explore the effects of regularization in direct data-driven predictive control (DDPC). In particular, we consider a nested bi-level DDPC,
$\mathrm{subject~to}\quad$ ${H}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (7a)
$\textrm{where}\quad{H}^{*}\in\arg\min_{{H}\in\mathcal{S}}\;J_{\mathrm{id}}({H},H_{\mathrm{d}}).$ (7b)
@@
The inner problem (7b) processes the offline data $H_{\mathrm{d}}$ (see (5)), which may be collected from non-deterministic LTI systems or nonlinear systems, and then the resulting predictor is used in the outer online control (7a). This bi-level formulation, often referred to as indirect DDPC, is modular and consists of two well-separated components: an inner identification layer and an outer model-based predictive control layer. Before discussing its connections and distinctions with existing schemes, let us further clarify the notation in (7). In the outer problem, the term $\|y\|_{Q}^{2}+\|u\|_{R}^{2}$ denotes the usual one $\sum_{k=t}^{t+N-1}\big(\|y(k)\|_{Q}^{2}+\|u(k)\|_{R}^{2}\big)$, and we have introduced a slack variable $\sigma_{y}$ with constraint $\Gamma$ and a regularization term $\|\sigma_{y}\|_{2}^{2}$ with $\lambda_{y}>0$. This ensures constraint feasibility despite noise, as used in [markovsky2021behavioral, berberich2020data]. In the inner problem, $J_{\mathrm{id}}(\cdot,\cdot)$ denotes a suitable loss function, and the constraint ${H}\in\mathcal{S}$ enforces prior data structures. We will clarify our choice of ${H}\in\mathcal{S}$ in Section 3.
Remark 1 (Comparison with [dorfler2022bridging]).
@@
The bi-level DDPC (7) is inspired by the work [dorfler2022bridging], which argues that “direct and regularized data-driven control can be derived as a convex relaxation of the indirect approach”. While adopting the same general framework, our work introduces two key distinctions. Conceptually, the bi-level DDPC (7) explicitly emphasizes the use of a non-parametric trajectory matrix as the predictive model, whereas [dorfler2022bridging] primarily considers a more general and abstract behavioral LTI setting in the inner problem. Technically, the analysis in [dorfler2022bridging] relies on arguments that require additional assumptions or clarifications, and some claims may not hold without these refinements. Also, the exact penalty arguments in [dorfler2022bridging] rely on [clarke1990optimization, Proposition 2.4.3], which requires adaptation in the DDPC context. Inspired by the notion of partial calmness, our proof is transparent and self-contained. We will clarify these points in more detail in Sections 3 and 4.
@@ 3.1 Behavior-based identification for predictive control
From the offline noisy data $H_{\mathrm{d}}$, we aim to get a new trajectory library $H^{*}$ for predictive control
${H}^{*}:=\textnormal{col}({U}_{\textnormal{P}}^{*},{Y}_{\textnormal{P}}^{*},{U}_{\textnormal{F}}^{*},{Y}_{\textnormal{F}}^{*}),\vskip-2.84526pt$
where each column of $H^{*}$ is a “purified” trajectory of the system. Let $\mathcal{W}_{L}$ (where $L=T_{\textnormal{ini}}+N$) denotes the space of all possible length-$L$ trajectories of an LTI system, i.e.,
@@
$H^{*}=\operatorname*{argmin}_{H}$ $J_{\mathrm{id}}({H},H_{\mathrm{d}})=\|H-H_{\mathrm{d}}\|_{F}^{2}$ (8)
$\mathrm{subject~to}$ $\mathrm{Im}(H)\subseteq\mathcal{W}_{L}.$
If the offline data $H_{\mathrm{d}}$ is noise-free and satisfies the persistent excitation, the fundamental lemma ensures that $\mathrm{Im}(H_{\mathrm{d}})=\mathcal{W}_{L}$. In this case, the optimal solution to (8) is $H^{*}={H}_{\mathrm{d}}$. In general, the constraint $\mathrm{Im}(H)\subseteq\mathcal{W}_{L}$ is too abstract and not tractable. For an explicit expression, we have several necessary conditions.
  • First, the trajectory matrix $H$ should be of low rank, i.e., we have $\mathrm{rank}\,H^{*}\leq mL+n$.
@@
    ${Y}_{\textnormal{F}}^{*}=K\ \textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}}),\vskip-5.69054pt$ (9)
@@
  • Third, the trajectory matrix $H$ should satisfy causality, i.e., the coefficient matrix $K$ in (9) should have a block partition $K=\begin{bmatrix}K_{\mathrm{p}}\;K_{\mathrm{f}}\end{bmatrix}$, and $K_{\mathrm{f}}\in\mathcal{L}$ encodes a block-lower triangle structure for causality.
  • Finally, we may also impose a Hankel structure on $H^{*}\in\mathcal{H}$ if each column of $H_{\textnormal{d}}$ is a one-step shifted trajectory. Representing the trajectory library with a Hankel structure is widely used in both theory [berberich2020data, berberich2022linear] and practical applications [elokda2021data, wang2023implementation] of DDPC.
@@
$({H}^{*},K^{*})\in\operatorname*{argmin}_{\tilde{H},K}\quad$ $\|\tilde{H}-H_{\mathrm{d}}\|_{F}^{2}$
@@
We thus consider (10) as our inner identification in (7b). Note that the requirements of linearity (10b) and causality (10c) are typically used in causal SPC [favoreel1999spc, fiedler2021relationship], and the requirement of low rank (10a) and Hankel structure (10d) are often employed in low-rank approximation [markovsky2016missing]. When the offline data $H_{\mathrm{d}}$ in (5) is collected from an LTI system (1) with noise-free data, the unique optimal solution to (10) is trivially ${H}^{*}=H_{\mathrm{d}}$. If $H_{\mathrm{d}}$ contains “variance” noises and/or “bias” errors, enforcing all constraints simultaneously makes the identification (10) difficult to solve due to its nonconvexity. We will demonstrate two strategies to address this challenge: 1) simplifying (10) via only considering a subset of the constraints (Section 4) and 2) approximately solving (10) via addressing each constraint sequentially (Section 5).
@@ 3.2 A direct approach via penalty methods
With the inner identification (10), the bi-level formulation (7) is an indirect DDPC. Using penalty methods, the bi-level structure can be transformed into a single-level problem. This is a standard idea in bi-level optimization [ye1997exact] and has been recently used to analyze direct data-driven control in [dorfler2022bridging]. We adopt a similar strategy here, but also highlight a subtle point that leads to some conceptual inaccuracies in [dorfler2022bridging].
@@
[…]
@@
Under a mild assumption (see e.g., [ruszczynski2011nonlinear, Theorem 6.6]), we can ensure that as the penalty parameter $\lambda\to\infty$, the optimal solution of (13) converges to an optimal solution of (11), i.e., the original bi-level problem (7). If we use a small penalty parameter $\lambda$, the single-level formulation (13) becomes a relaxation of (7), i.e., the optimal value of (13) is smaller than that in (7).
The reasoning above critically relies on the conceptual optimality constraint $H\in\mathcal{C}_{\mathrm{opt}}$, associated with the inner problem (7b). Dropping any constraint in (7b) results in a different optimality set, meaning that the corresponding penalized single-level problem may no longer offer a valid relaxation. This subtle but important point may be easily overlooked; for example, some relaxation statements in [dorfler2022bridging] are not fully accurate without further clarification. We consider a simple example below.
@@
It can be verified that when $\lambda\geq\frac{1}{4}$ (see Fig. 2 for illustration), (15) has the same optimal solution as (14). If $\lambda<\frac{1}{4}$, the optimal value of (15) is a lower bound of (14). If we relax the constraint (14c) as $y\leq 4$, then the inner problem has a different optimal solution $y^{*}=0$. In this case, the outer problem in (14a) becomes infeasible. This illustrates that dropping constraints from the inner problem does not necessarily reduce the optimal value of the outer problem. This nuance was overlooked in [dorfler2022bridging].